OpenStudy (anonymous):

Write the solution for the following problem as an ordered pair: The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Be sure to show all work.

OpenStudy (anonymous):

46=perimeter 46=x+x+y y=x+4 x+x+x+4=46 can you go from here?